### v2

# FSLH: Flexible Mechanized Speculative Load Hardening FSLH: Flexible Mechanized Speculative Load Hardening

Abstract
The Spectre speculative side-channel attacks pose formidable threats for computer system security. Research has shown that cryptographic constant-time code can be efficiently protected against Spectre v1 using a selective variant of Speculative Load Hardening (SLH). SLH was, however, not strong enough for protecting non-cryptographic code, leading to the introduction of Ultimate SLH, which provides protection for arbitrary programs, but has too large overhead for general use, since it conservatively assumes that all data is secret. In this paper we introduce a flexible SLH notion that achieves the best of both worlds by formally generalizing both Selective and Ultimate SLH. We give a suitable security definition for such transformations protecting arbitrary programs: any transformed program running with speculation should not leak more than what the source program leaks sequentially. We formally prove using the Rocq prover that two flexible SLH variants enforce this relative security guarantee. As easy corollaries we also obtain that Ultimate SLH enforces our relative security notion, and also that the selective variants of value SLH and address SLH enforce speculative constant-time security.
Keywords:
side-channel attacks, speculative execution, Spectre, secure compilation, speculative load hardening, speculative constant time, relative security, formal verification, Rocq, CoqI Introduction
Speculative side-channel attacks such as Spectre pose formidable threats for the security of computer systems [8, 19]. For instance, in typical Spectre v1 attacks [18], misspeculated array bounds checks cause out-of-bounds memory accesses to inadvertently load and reveal secrets via timing variations. SLH [9] is a software countermeasure against such attacks, originally proposed and implemented by LLVM, that dynamically tracks whether execution is in a mispredicted branch using a misspeculation flag register. This misspeculation flag is used as a mask to erase the value of any misspeculated loads, in a variant Zhang et al. [26] denote as vSLH, for value SLH. Another variant supported by LLVM is to mask the address of any misspeculated loads, which also following Zhang et al., we denote as aSLH, for address SLH.
It is, however, challenging to build software protections that are both efficient and that provide formal end-to-end security guarantees against precisely specified, speculative side-channel attacker models [12]. Cryptography researchers are leading the way in this space, with defenses such as selective vSLH efficiently achieving speculative constant-time guarantees against Spectre v1 for cryptographic code with typical overheads under 1% [22, 23]. This work is, however, specialized to only cryptographic code and often also to domain-specific languages for cryptography, such as Jasmin [2].
It is more difficult to properly protect arbitrary programs written in general-purpose languages, which in particular do not obey the cryptographic constant-time discipline [1, 10]. SLH was not strong enough for protecting such non-cryptographic code [20], leading to the introduction of Ultimate SLH [20, 26], which uses the misspeculation flag to mask not only the values loaded from memory, but also all branch conditions, all memory addresses, the operands of all non-constant time operations, etc. While this should in principle be strong enough to achieve a relative security notion [12, 26, 20], it also brings 150% overhead on the SPEC benchmarks [26], which seems unacceptable for many practical scenarios.
In this paper we introduce FSLH, a flexible SLH notion that achieves the best of both worlds by generalizing both Selective and Ultimate SLH. Like Selective SLH, FSLH keeps track of which program inputs are secret and which ones not and only protects those memory operations that could potentially leak secret inputs when ran speculatively. Like Ultimate SLH, FSLH also provides protection for non-cryptographic code that does not respect the constant-time discipline, but does this by only masking those branch conditions and memory addresses that are potentially influenced by secrets.
Contributions
We introduce FSLH, a Flexible SLH notion generalizing both Selective SLH and Ultimate SLH. This applies to both vSLH and aSLH, resulting in two variants: Flexible value SLH (FvSLH) and Flexible address SLH (FaSLH).
We prove formally in the Rocq prover111The Rocq interactive theorem prover was previously known as Coq. that FvSLH and FaSLH can each be instantiated to recover both the corresponding Selective SLH variant as well as Ultimate SLH. The connection to Selective SLH guarantees that cryptographic constant-time code pays no extra performance penalty under Flexible SLH compared to Selective SLH. The connection to Ultimate SLH states that if all data is labeled as secret then the program receives the same protection as with Ultimate SLH.
We give a suitable relative security definition for transformations protecting arbitrary programs, like FSLH and Ultimate SLH: any transformed program running with speculation should not leak more than what the source program leaks sequentially (so without speculation).
We prove in Rocq that our two flexible SLH variants, FvSLH and FaSLH, enforce this relative security notion.
Other easy corollaries we obtain in Rocq are that Selective vSLH and Selective aSLH enforce speculative constant-time security, whereas previously only Selective vSLH was proved secure and only on paper [22]. In fact, for Selective aSLH we are the first to even give a formal definition, which is nontrivial for memory stores.
Outline
§ II presents required background, after which § III introduces our key ideas on FaSLH. § IV defines our simple imperative language and its sequential and speculative semantics. § V describes the proofs of our formal results for FaSLH, while § VI quickly presents the analogous results for FvSLH. § VII describes related work and § VIII concludes with future work.
Artifact
All the results of this paper have been fully formalized in
the Rocq proof assistant and are available at https://nce.mpi-sp.org/index.php/s/nggS2PMEWa5gynJ
The Rocq development leading to our main theorems has 2600 lines of code.
II Background
II-A Speculative Execution Attacks
Modern processors implement a variety of hardware mechanisms such are caches to get software to run faster. Caches are part of the internal microarchitectural state of the CPU: they improve performance, but are not exposed to the programmer through the instruction set architecture (ISA) and thus have no bearing on the result of any computations. Yet their effects on performance can be detected by measuring the execution time of a program, which may leak information about its data. Timing side-channel attacks exploiting microarchitectural leakage are a common concern for cryptographic code and other secret-manipulating programs.
Speculative execution is another technique designed to increase performance by executing instructions that are likely, even if not certain, to be needed by a program. For example, the condition of a branch instruction could depend on slow memory accesses or the results of instructions that have not finished executing. Instead of waiting, a processor can use a branch predictor to guess an outcome for the condition and immediately start executing instructions from the chosen branch. When the result of the condition is later known, the processor commits the effects of the speculative path if the guess was correct, or reverts them and starts executing the correct path if the guess was wrong. This rollback ensures that misspeculation has no effect on ISA state, but does not undo changes to the microarchitectural state, like the caches.
This opens the door to speculative leaks via timing side channels that would not exist if there was no misspeculation. Even more, an attacker can train the predictors to misspeculate (directly if collocated, or even indirectly by causing the program’s own code to be cleverly invoked [21]), which is the essence of Spectre attacks. A Spectre v1 attack [18] (aka Spectre-PHT [8]) targets the branch predictor to steer a target program down certain speculative paths that create timing leakage. This is the classic speculative execution attack:
Listing 1 (Spectre v1 gadget).
| ifi < a1_sizethen | ||
This code indexes a secret array over an untrusted input i. A conditional validates that the input is in the right range, and only in this case accesses at that position. It then uses the result to index a read from a second array . The problem is that prior to invoking this code an attacker can train the branch predictor to guess that the then branch will be taken. The attacker can then run the program with an out-of-bounds input i that actually points at secret data, and although the then branch should not be taken, the processor can start running the instructions on the then branch before the actual value of the branch condition is resolved. The first array access loads out-of-bounds secret data into j, and the second access loads from at an address that depends on that j. While the processor will course-correct as soon as it detects its misspeculation, the access to leaves a footprint in the cache, so the attacker can measure timing to infer the value of j, which contains speculatively loaded secret data.
II-B Speculative Load Hardening
Speculative Load Hardening [9] (SLH) is a family of software-based countermeasures against Spectre v1 attacks originally proposed and implemented by LLVM. SLH is a program transformation that combines two ingredients:
1.
a mechanism to keep track of misspeculation, and
2.
a means of preventing speculative leaks.
The first is shared by all SLH variants. The core transformation from Figure 1 shows how to maintain a misspeculation flag tracking whether any branch was mispredicted along the current control path at any given time. This is possible because all modern architectures have support for implementing control flow conditions using branchless and unpredicted instructions, e.g., conditional moves in x86 assembly.
We can easily implement such a flag in a language that uses those instructions to provide a constant-time conditional, like the operator in Figure 1, which evaluates if is true, and if it is false. We assume the variable b is reserved for the flag and unused by programs, and we ensure that its value is 1 if the processor is misspeculating, and 0 if it is not. The transformation applies recursively to the structure of commands. Ignoring for now the helpers that parameterize the scheme (, and ), the interesting cases are the two branching commands: conditionals and loops.
In both cases there are two paths: one where the branch condition is true and one where it is false. For instance, if the branch condition is true, the command in the then branch (resp. in the loop body) works assuming that is true. We use a constant-time conditional to evaluate the same condition and determine if the processor is misspeculating. If it is not, the condition will indeed evaluate to true, and we will leave the speculation flag unchanged (i.e., no new misspeculation has occurred). If it is misspeculating, the condition will evaluate to false, and we will set the flag to 1. This allows the program to detect misspeculation and protect itself accordingly.
The second ingredient of the transformation is using the misspeculation flag to prevent speculative leaks. Figure 1 provides a generic template for aSLH, where the indexes of array loads and stores, as well as the branch conditions can be sanitized in different ways, resulting in different variants, each with different trade-offs between their degree of protection against speculative side channels and their performance cost.
Let’s first consider the standard aSLH transformation, which leaves branch conditions unchanged, i.e., , and protects all array accesses based on the speculation flag. If b contains the speculation flag and i is the index we intend to access, we can use the constant-time conditional expression b=1 ? 0 : i to protect it: when the processor is misspeculating (as indicated by the set flag b=1) the index is zeroed and the access protected; when the processor does not misspeculate, the access is left unchanged. That is, .
Applying aSLH to 1 we obtain the following code:
Listing 2 (Spectre v1 gadget protected with aSLH).
| ifi < a1_sizethen | ||
| else | ||
| b:=(i < a1_size)?1:b |
A Spectre v1 attacker can still force the execution of the then branch when i is out of bounds, but this will be detected and b will be updated to reflect it. Then the first array access will become and will no longer be out of bounds, and the attacker will be unable to exploit it to load secrets.
II-C Cryptographic Code Security
Cryptographic implementations are common targets of timing side-channel attacks aimed at extracting keys, even remotely over the Internet [7]. To mitigate these dangers, cryptographic engineers have developed secure programming disciplines, notably cryptographic constant time (CCT), now considered standard practice in the field [1, 10, 15, 4]. CCT protects secrets from timing side-channel attacks arising from secret-dependent branches and memory accesses. For this, it imposes two requirements on the programmer:
1.
all program inputs must be identified as public or secret,
2.
control flow and memory addresses that depend on secret inputs are not allowed.222 Also variable-time arithmetic operations are not allowed on secrets, but these are not relevant for the simple setting of this paper.
Researchers have designed security analyses ensuring that these CCT conditions are met for real-world code [1, 15, 4]. Even if the attacker can directly observe the branches that are taken and the memory addresses that are accessed, a CCT program does not leak any secret information to the attacker. Prior work has shown that this implies security against timing side-channel attacks exploiting data and instruction caches [4].
Definition 1 (Cryptographic Constant-Time Security).
A program satisfies CCT security w.r.t. public variables if for any public-equivalent states (), when the program executes to completion on both states (i.e., a small-step semantics fully evaluates the program from each state), then both runs produce the same sequence of branching decisions and memory accesses observable by an attacker ().
In the speculative world, an attacker can invalidate these guarantees. By influencing the microarchitectural state of the hardware, it can force a program to speculatively execute control paths that are impossible in the standard, sequential semantics, and leak secret data through the timing side channels CCT is supposed to close [11, 5, 24, 22]. To reason about these threats, we need a more informative semantics that abstracts the speculative capabilities of an attacker [22]. This speculative semantics ( below) adds rules that steer execution along wrong branches and out-of-bounds memory accesses, triggered by directives issued by the attacker. The speculative semantics also adds a boolean flag to the state to indicate whether the processor is misspeculating. We define speculative constant-time (SCT) security below in terms of a notion of speculative observationally equivalence:
Definition 2 (Speculative observational equivalence).
Two speculative configurations and are observationally equivalent w.r.t. the speculative semantics, written iff
Compared to SCT definitions from the literature [11, 14, 24] we are looking at all execution prefixes that do not necessarily finish in a final state (the skip in 1). This choice gives us stronger guarantees in the definitions below, for instance making our SCT definition termination-sensitive. We can directly compare observations and for equality here, since the two speculative executions take identical directions , which in our setting implies that .
Definition 3 (Speculative Constant-Time Security).
A program satisfies SCT security if for any public-equivalent initial states and starting before mispeculation ()
Example 1.
It is easy to see that 1 is CCT if the variables are initially public. Even so, speculatively it leaks secrets, as discussed in § II-A. Formally, this program is not SCT secure, since for an out-of-bounds i, the attacker can use a first directive in to force the program to misspeculate on the then branch, and then a second directive to ask for a secret in memory instead of the out-of-bounds value of . The secret is loaded in j, and then made visible to the attacker through the address of the second load operation, which is observed to have different values in and . On the other hand, after aSLH protection (2) this program is SCT secure. In fact, any CCT program hardened with aSLH satisfies SCT security, and as we will see in the next subsection, a selective variant of aSLH is enough for enforcing this.
III Key Ideas
III-A Selective aSLH
Selective SLH is a more efficient variant of SLH that “only masks values speculatively loaded into publicly-typed variables” [22], exploiting the security levels present in CCT programs to minimize the number of masking operations while protecting secrets from speculative attackers. The original proposal uses a CCT type system to only mask the values of loads into public variables (i.e., the SvSLH variant we discuss later in § VI). Yet the same optimization is also applicable to aSLH leading to Selective address SLH (SaSLH), which we show in 2(a), and which as far as we know, has not been studied formally before. The masking of loads in uses a public map P of variables to their boolean security levels (public or secret ) and functions and computing the security levels of arithmetic and boolean expressions.
One crucial observation we make in this paper is that for achieving security for SaSLH it does not suffice to only mask the addresses of loads, as shown by the following counterexample we discovered during the formal proofs:
Listing 3 (Leakage through unprotected stores).
| ifi < a1_sizethen | ||
Once again for an invalid i the attacker can force the program to execute the then branch. The first instruction stores a secret value, key, normally into a secret array secrets, but with the out-of-bounds index this actually stores the secret at position 0 of a public array a. A subsequent and seemingly innocuous in-bounds load from a[0], where the secret key was already exfiltrated, then leaks its value to the attacker via a subsequent timing observation (). Since the a[0] load is from (the in-bounds) index 0, even though speculation is detected, the aSLH masking of this load will be a no-op and thus alone will not prevent this counterexample.
The solution is to define in 2(a) to mask the indexes of array stores when writing secret values. This extra masking causes the store operation in the example above to write the key in position 0 of secrets, leaving a unchanged and thus preventing the attack. With this addition we were able to formally prove speculative constant-time security:
Theorem 1.
SaSLH enforces SCT security for CCT programs.
III-B Defining Relative Security for Ultimate SLH
Most programs do not satisfy the CCT policy, so lightweight SLH transformations like the two aSLH variants above are not enough to enforce security for arbitrary programs.
Listing 4 (Leakage through sequentially unreachable code).
333This counterexample was shared with us by Gilles Barthe.The code in the outer then branch will never be sequentially executed, but a speculative attacker can force the execution of this code. The inner branch condition reveals information about the secret variable, and no amount of masking of memory accesses can prevent this. Since it branches on secret, this code is not CCT, but this is the kind code we still want to protect from Spectre v1 attackers. To achieve this all sources of speculative leakage observations must be masked, including branch conditions, not only the addresses of memory loads and stores.
Ultimate SLH (USLH) [20, 26] is a SLH variant that offers protection to arbitrary programs by exhaustively masking these three sources of leakage, as shown in 2(c). Like vanilla aSLH, it masks all addresses of memory loads and stores. Additionally, it also protects branch conditions by masking them with the misspeculation flag. This causes branches to default to the false case during misspeculation, preventing them from leaking information about the original branch conditions. This prevents the leakage in footnote 3.
Yet what is a suitable formal security definition for transformations protecting arbitrary programs, like USLH? Because the arbitrary source program does not follow the CCT discipline it could leak some data sequentially, and the transformations we look at here do not try to prevent these sequential leaks. Instead, defenses like USLH enforce that the hardened program does not leak any more information speculatively than the source program leaks sequentially. That is, if a sequential attacker against the source program cannot tell apart two input states, the transformation ensures that a speculative attacker cannot tell them apart either. For formalizing this relative security notion we first also define a notion of observationally equivalent configurations in the sequential semantics:
Definition 4 (Sequential observational equivalence).
Two source configurations and are observationally equivalent in the sequential semantics, written , iff
Looking at all execution prefixes above can result in observation sequences of different lengths, so we require that one sequence is a prefix of the other ().
Definition 5 (Relative Security).
A program transformation satisfies relative security if for all programs run from two arbitrary initial states and we have
Looking at all execution prefixes in the sequential observational equivalence premise of this definition is not a choice, but is forced on us by the fact that a program that sequentially loops forever or ends with an error (i.e., gets stuck before reaching skip) can be forced by attacker directions to successfully terminate speculatively. So if we were to only look at successfully terminating sequential executions (like in 1) in the premise, we would simply not have any information about the sequentially nonterminating or erroneous executions, and we would thus not be able to prove (even a big-step version of) the conclusion when such programs only successfully terminate speculatively.
Beyond these details, the key idea is to relate the security of the transformed program executed speculatively to the security of the source program executed sequentially. This provides stronger guarantees than in previous work [20, 26], as discussed in § VII. This definition also suitably captures the security guarantees offered by USLH, and later FSLH.
Theorem 2.
USLH enforces relative security for all programs.
III-C The Best of Both Worlds: Flexible aSLH
SaSLH and USLH can be seen as falling on two rather different points on the spectrum of Spectre v1 protections:
SaSLH exploits information from the developer identifying which inputs are secret and which ones are not to implement an efficient countermeasure that under the very strong requirement that the source program is CCT guarantees no leakage of secret inputs whatsoever.
USLH pays a much higher efficiency price to protect arbitrary programs and obtain a stronger relative security guarantee that includes programs that are not CCT.
In this paper we propose Flexible address SLH (FaSLH), a new hybrid design shown in 2(b) that combines the desirable features of both SaSLH and USLH:
Like SaSLH, FaSLH also exploits information from the developer identifying which program inputs are secret and which are not, and only offers speculative protections to data that could depend on secret inputs. Unlike SaSLH though, FaSLH does not require the strict CCT discipline.
Like USLH, FaSLH aims to offer relative security to arbitrary programs, which in our current formalization are only well-typed with respect to a standard information flow control (IFC) type system tracking explicit and implicit flows [25] (see § IV-C).444As discussed in § VIII, replacing this IFC type system with a flow-sensitive IFC analysis accepting all programs should be possible in the future. Unlike USLH though, FaSLH uses the information about which values may have been influenced by secret inputs and which ones not to only selectively apply SLH protections, as detailed below.
As shown in 2(b), the treatment of branch conditions is a hybrid between USLH (whose unique contributions are given in red) and SaSLH (in blue), with some new logic (in purple). When the branch condition is a secret expression, it behaves like USLH, using the misspeculation flag to mask the branch condition. When the branch condition is public, it behaves like SaSLH, and leaves the branch condition unchanged.
For array loads, while USLH always masks the indices, SaSLH only masks indices for loads that sequentially involve public data. This is, however, not enough for relative security of non-CCT programs: since speculative stores leak the index, FaSLH also has to mask secret indices to prevent the following speculative leak (similar to footnote 3 above).
Listing 5 (Leakage through sequentially unreachable load).
This is similar for array stores, while USLH masks indices for all stores, SaSLH only masks indices for stores if the expression being stored is secret. This is again not enough for relative security of non-CCT programs: since stores leak the index, FaSLH also masks secret indices of stores.
Listing 6 (Leakage through sequentially unreachable store).
We have proved in Rocq that the masking done by FaSLH is enough to enforce the following variant of 5:
Theorem 3.
FaSLH enforces relative security for all IFC-well-typed programs and for all public-equivalent initial states.
The extra premise that the initial states are public-equivalent intuitively captures the requirement that the developer has properly identified the secret inputs they want to try to protect, since all other inputs will not be protected by FaSLH. Perhaps a bit counterintuitively, because of the other, semantic equivalence premise from 5, an input labeled as secret is protected speculatively by FaSLH only if it does not leak sequentially. This is the most one can achieve for non-CCT programs, without heavier defenses enforcing CCT [10].
Finally, we proved in Rocq that FaSLH is a generalization of SaSLH and USLH. The former means that CCT code pays no extra performance penalty under FaSLH compared to SaSLH:
Theorem 4 (Connection between FaSLH and SaSLH).
For any CCT program and public variable map P we have
On the other hand, if all data is labeled as secret then the program receives exactly the same protection as with USLH:
Theorem 5 (Connection between FaSLH and USLH).
If all variables (both scalars and arrays) are labeled as secret (i.e., we use public map ), then for any program we have
IV Definitions
IV-A Language Syntax and Sequential Semantics
Throughout the paper we use a simple while language with arrays we call AWhile. Its syntax is shown in Figure 3. A first layer comprises pure arithmetic expressions (aexp) on natural numbers (which we also write , or if they compute array indexes), and boolean expressions (bexp) (also written ). The second layer of commands (com) includes standard imperative constructs, as well as array accesses.
| number | |||||
| scalar variable | |||||
| arithmetic operator | |||||
| constant-time conditional | |||||
| boolean | |||||
| arithmetic comparison | |||||
| boolean operator | |||||
| skip | do nothing | ||||
| assignment | |||||
| sequence | |||||
| conditional | |||||
| loop | |||||
| read from array | |||||
| write to array |
Program states are divided into scalar variables and arrays . A register file assigns values to scalar variables. This mapping is used to evaluate arithmetic and boolean expressions involving scalar variables in the usual way, written . To update the value of a variable in the mapping we write . The second half of the program state is a main memory that stores the arrays. The mapping assigns a fixed size to each array, and defines a lookup function to fetch the value of an array at a given valid index, i.e., in the range . To update the value of an array at a valid index we write . The contents of an array cannot be used directly in arithmetic or boolean expressions, and can only be transferred to and from scalar variables through the read and write commands.
We define a standard small-step operational semantics for AWhile (for details, see § IV-B later and also § A). States are triples composed of the next command to execute, and the scalar and array states. We write the step or if we do not care about the state after stepping. A step produces an optional observation that represents the passive capabilities of an attacker to obtain information about the program via side channels. We use the standard model for sequential side-channel attackers, which observes the control flow followed by program ( for a branch condition that evaluates to a boolean ) and the accessed memory addresses ( and resp. for array reads and writes at a[i]). If a step does not produce an observation, we write instead of a concrete event. Multi-step execution is the reflexive and transitive closure of the step relation, written , where is a trace of events. In our formalization silent steps leave no mark in the trace. Full executions are multi-step executions of the form .
It is important to note that evaluating arithmetic and boolean expressions does not produce any observable event, even though there is a strong resemblance between the conditional expression and the conditional command . Internally, this imposes the use of branchless and unpredicted logic to implement , and is critical to the correctness of our SLH countermeasures.
Example 2.
Let us revisit 1 using this formal model. There are only two ways to fully evaluate that code under the sequential semantics. Without loss of generality, suppose contains the 4-element array and another large array , say of size 1000. Suppose also that . Then, depending on the prospective index on given by i:
1.
If the access is valid, e.g., , the conditional produces an observation , and afterwards the then branch executes, emitting two observations: and (the largest element in is 7, so all accesses on based on the contents of will be valid).
2.
If the access is not valid, e.g., , the conditional produces an observation , and the program terminates without any additional observations.
We can see that under these conditions the code is sequentially secure, i.e., out of bounds accesses cannot occur. If we have a pair of states that agree on the attacker-supplied value of i (and the program data, as described above), then the premise of 5 is also satisfied.
IV-B Speculative Semantics
The observations produced by executions in the sequential semantics allow us to reason about leakage that follows necessarily from the intended control flow and memory access patterns of the program. This is enough for properties like CCT in the standard leakage model, where the attacker is passive. However, a speculative attacker is able to actively influence the program by causing it to veer outside the paths prescribed by the sequential semantics, and these transient paths empower the attacker to make additional observations about the state of the program. To model these more powerful attackers, we use a speculative semantics [22] that reflects these active capabilities, shown in Figure 4. (This is a strict extension of the sequential semantics, which we recover by removing the additions highlighted in red.)
The small-step relation is a generalization of the sequential relation that extends it in two ways. The first change is the addition of an optional speculation directive that abstracts the active capabilities of the attacker to steer an execution down a specific speculative path. Each observation-producing rule in the sequential semantics is split into two speculative variants. The first variant behaves exactly like its sequential counterpart, and does so when the attacker gives the green light with the directive step. The second variant misspeculates, executing a branch that should not be taken (when issuing directive force for conditionals), or performing an arbitrary array access (when issuing directives and resp. for reads and writes). Note the initial speculation flag in Spec_Read_Force and Spec_Write_Force: the attacker can only cause misspeculation on array accesses after having forced a branch misprediction, and only when it manages to force an out-of-bounds access. In this case, we conservatively give the attacker choice over the resulting access. This overapproximate the speculative capabilities of a Spectre v1 attacker. Again, we write when no directive is needed.
The second change is the extension of program states with a fourth component, a boolean flag that indicates whether the program has diverged from its sequential execution. The flag is set to at the start of the execution, and updated to when the attacker forces a branch misprediction.
Again, we write for multi-step execution, where is the sequence of speculative directives selected by the attacker. Because our security definitions quantify over all possible execution paths, the speculative semantics does not need to model implementation details like finite speculation windows or rollback [22], as also discussed in § VII.
Example 3.
Using the sequential and speculative semantics for AWhile, we can show that 1 without protections is not relative secure, i.e., does not satisfy 5.
Assume the scalar variables i and a1_size and the arrays and are public, and all other data is secret. As we have seen in 2, the program satisfies 1, and therefore for any pair of public-equivalent states also 4. But this does not mean that two executions from related states with the same directives will yield identical observations. Let us return to the second sample scenario in the last example, where in an attempt to access out of bounds. Suppose that and are identical except for the fact that contains a 1-element array , but in we have . The attacker can falsify the property as follows:
1.
Issue a directive force to speculatively begin executing the instructions on the then branch, even though the condition is actually false. This produces an observation on both runs and initiates a misspeculated path.
2.
The first read is an out of bounds access, so the attacker can choose a directive . This produces an observation on both runs. At this point the scalar states differ: j holds the secret 42 in and 43 in .
3.
The second read is allowed to proceed normally using step, but it is too late: this produces the observation on the first run and on the second, which reveals the secret and violates the property.
IV-C Simple IFC Type System
As mentioned in § III-C, our current implementation of flexible SLH in Rocq make use of a simple IFC type system à la Volpano-Smith [25] that tracks explicit and implicit flows in AWhile programs (Figure 5). For this we make use of a pair of maps that assign security labels to scalar variables () and to array variables (), respectively. We consider the two-point lattice of booleans levels, with for public and for public. Since public may flow into the secret level we take . We lift the map of public variables P to a pair of functions and that compute the security levels of arithmetic and boolean expressions in the usual way. We define a standard public-equivalence relation that relates two scalar states iff they agree on the public variables according to P agree, and similarly iff they agree on the sizes and contents of the public arrays according to PA.
We define the standard typing judgment using the rules from Figure 5. This will ensure that command , run from two public-equivalent states is secure, in the sense that it never produces any observations that leak secret data.
It will be interesting below that this type system generalizes the type systems used for CCT (e.g., in § II-C). Figure 5 highlights the extensions of our type system w.r.t. the corresponding CCT type system in red: implicit flows are tracked through the PC label, the labels of branching conditions and memory addresses are no longer required to be public, and joins are generalized to account for these new labels and flows.
V Formal Results
The main goal of this section is to prove that FaSLH enforces relative security (5), and from there to derive security results for other SLH variants that are special cases of the more general FaSLH.
V-A Ideal Semantics
Similarly to [22], a technical device used by the security proofs is an auxiliary semantics that refines the speculative semantics from Figure 4. This ideal semantics introduces new restrictions which reflect the idealized behavior of programs that are hardened against speculative leaks by FaSLH.
Again, we define the semantics as a small-step relation and write . The labeling map P is constant throughout, so we elide it and the turnstile from the relation, although we may refer to it inside premises. The speculative and the ideal semantics share the same collection of rules. Figure 6 shows the subset of rules that change from the speculative semantics, with additions highlighted in red.
The changed rules are exactly the rules that generate observations, and the changes are exclusively additions to premises. The two rules for conditionals simply mask the branch condition so that it defaults to the else branch in some cases. Similarly, the step rules for array loads and stores add conditions under which the index (i.e., the address) is masked instead of evaluated, and in these cases defaults to the first element of the array. The misspeculating rules on array loads and stores only apply under certain safe scenarios, and get stuck when these are not met. Finally, multi-step executions and ideal observational equivalence are defined in precisely the same way as their speculative counterparts.
Example 4.
The ideal semantics only allows misspeculating read on an array to proceed if, in addition to forcing an out-of-bounds access as in the speculative semantics, this access would load from a public address to a secret variable.
A read operation on an array is only allowed to misspeculate if the index is public and if the result is stored to a secret variable. This comes from the fact that allowing the operation leaks the index, so it has to be public or masked. Moreover, misspeculating allows the attacker to read from any array, even if the index is public. For example, in 1, a read out of bounds after misspeculating in the branch allows the attacker to store basically any part of the memory in j, even when i is public information.
The way misspeculation on arrays is prevented is by masking the index to 0 when the speculation flag is set to . This ensures that reads always happen inside of bounds, so no misspeculation is possible.
The design of the ideal semantics restricts the speculative semantics in such a way that it yields additional information for use in the proofs that follow.
V-B Key Theorems
There are several key lemmas used in our proof of Theorem 3. The first of these is a backwards compiler correctness (BCC) result showing that the program hardened by FaSLH executed with the unrestricted speculative semantics behaves like the original source program executed with the ideal semantics, i.e., the FaSLH translation correctly implements the restrictions of the ideal semantics. Despite their parallels, establishing the links between the two is technically subtle.
Lemma 1 (Backwards compiler correctness).
Explanation and proof outline.
This lemma connects a “target run” of the hardened command in (2) to a “source run” of the original command in (3). The initial configurations of both runs are identical except for the translation function; the memories and the speculation flag in the final states are also identical. Additionally, it concludes in (4) that terminating target runs correspond to terminating source runs, and at this point the values of speculation flag maintained by the program and the semantics coincide.
The result holds under a number of side conditions detailed in (1). On the one hand, the original program does not use the variable reserved for the speculation flag (this is also why b remains unchanged throughout the source run), and its initial value is the same as the semantic flag ; the function simply produces the natural encoding of the input boolean. On the other hand, we require all arrays to be non-empty, so that when we mask an index to 0 this does not result in an out-of-bound access that could be misspeculated upon.
The proof proceeds by induction on the number of steps in the target run (2), followed by a case analysis on the command. Note however that the number of steps cannot be inferred from the size of either the observations or the directives due to the presence of silent steps. Instead, we perform a strong induction on , where we add a measure of the program defined simply as the number of its constructors. This compound measure effectively bounds the number of steps taken by the speculative semantics in the target run. Loops are handled by the size of the trace of observations, as each iteration produces at least one observation. ∎
The standard formulation of BCC, without the additional conjuncts in (4), is a trivial corollary.
An important property of the ideal semantics used in the lemmas that follow allows us to relate the results of a pair of runs from related states when they are given the same directives and produce the same observations. Here we show the single-step version, but we have generalized this to multi-steps by ensuring that and cannot reduce without producing an observation.
Lemma 2 (Noninterference of ).
Proof sketch.
By induction on the first evaluation judgment and inversion on the second one. ∎
Another key technical result relates the behaviors of well-typed programs running from public-equivalent but misspeculating states, still in the ideal semantics [26].
Lemma 3 (Unwinding of ideal misspeculated executions).
Proof.
The proof works by induction on one of the speculative executions exposed after unfolding (2). The main auxiliary lemma is a stepwise version of this same statement that takes one step in each execution with the same directive and well-typed command, and again from public-equivalent states, and shows that the two steps produce the same observation and reduction of the command. This follows from the initial hypothesis (1), 2 and the fact that the step relation preserves well-typedness. ∎
The final piece of the puzzle requires showing that the ideal semantics satisfies relative security by itself, without the need for any hardening.
Lemma 4 ( ensures relative security).
Proof.
Unfolding the definition in (3) exposes the two ideal executions together with their shared directives and their observations and . The goal of the theorem is to establish their equality based on the equivalence of observations in the sequential semantics, given by (2), and the well-typedness and public-equivalence of the initial states in (1).
If the attacker never forces misspeculation, i.e., contains exclusively step directives, the goal follows directly because ideal executions without misspeculation preserve the property in (2), that is, they are identical to sequential executions.
If the attacker does force the program misspeculate at some point during the executions, the list of directives is necessarily of the form : a prefix without misspeculation, followed by a directive force that initiates the misspeculation on a branch, and then by an arbitrary suffix of directives. This neatly divides the proof into two: the prefix phase is identical to the first case up to the point where misspeculation begins, and the two ideal runs have reduced to the same command. 2 shows that the states reached after the pivot directive force are equivalent. At this point we apply 3 with the suffix to conclude the proof. ∎
Now we are ready to prove our main theorem.
Theorem 3 (FaSLH enforces relative security).
Proof.
The proof of relative security depends on some general assumptions in (1) about the original program not using the reserved variable for the misspeculation flag, and the flag being initialized to 0 in both memories. Together with these, there is well-typedness and public-equivalence in (3) and the requirement on array sizes in (2).
The other two top-level security results for other variants of SLH are simple corollaries of this theorem. Observe that the assumptions of Theorem 3 can be divided into general-purpose assumptions (in black), assumptions that are closely related to the security of SaSLH in Theorem 1 (in blue), and those related to the security of USLH in Theorem 2 (in red).
Theorem 1 (SaSLH enforces SCT security).
Proof.
From Theorem 4, SaSLH is identical to FaSLH for a fixed P. Together with the use of the more restrictive CCT type system instead of the general IFC type system, this imposes that all observations are based on public values, and we can deduce that all sequential executions from public-equivalent states produce the same observations. After deriving the missing assumption, we conclude by applying Theorem 3. ∎
Theorem 2 (USLH enforces relative security).
VI Flexible Value SLH
In the last section we formalized address SLH, which protects programs from speculative leakage by masking (some of) the addresses that produce memory accesses. For load operations specifically, an alternative is to target the “output” of the operations, i.e., prevent the values read from memory from leaking to the attacker. This has produced countermeasures like Selective value SLH.
We can adapt Flexible SLH to use value SLH countermeasures while generalizing existing schemes and offering similar protections against speculative attackers. Our general template for vSLH transformations is almost identical to the aSLH master recipe. Only the rule for array reads changes as shown in Figure 7. In aSLH, array reads were parameterized by an index translation function that was in charge of masking the address as needed. In vSLH, the translation of reads is split into two cases: one where the loaded value is immediately masked using the misspeculation flag, and one where it is not; the rule is parameterized by a value check function that controls the case analysis.
Observe that the vSLH template actually uses an index translation function in the case where the read value is not masked, as well as the index masking function for array writes, like aSLH did. On the one hand, SvSLH, gets its security exclusively from masking values that are loaded to public variables, and never protects indexes, as seen in 8(a). On the other hand, as shown in 8(b), FvSLH needs to fall back on masking certain addresses if it is to achieve relative security. Moreover, USLH can also be derived from FvSLH, and simply ignores its value masking facilities by having and falling back on index masking for all its protections.
Example 5.
Consider again 3. Both aSLH and vSLH prevent the public variable x from containing the secret value of key, but their strategies differ. In the aSLH setting, it is the out-of-bounds store that is prevented. On the other hand, vSLH does not protect the write operation as long as the index i is public. The attacker is indeed able to load the secret key into the public array a, which breaks the public-equivalence between arrays. In order to maintain security, the subsequent load value on x will be masked 0, and a[0] containing a secret does not cause the program to leak information.
Using these new definitions we can state and prove that FvSLH enforces relative security and SvSLH enforces SCT security. The high-level structure of the security proofs closely mirrors the development in § V. FvSLH uses a new version of the ideal semantics that reflects the changes in its behavior and supplies the proofs with relevant information. By a slight abuse of notation, we will write for the new semantics in this section and elide the public variables and turnstile as those remain constant throughout.
The new ideal semantics is in large part identical to the ideal semantics for FaSLH in Figure 6; the only changes take place in the premises of the rules for array reads and writes, which are given in Figure 9. Changes between the two versions are highlighted in red; premises that disappear in the new version are also crossed out. Both rules for array reads now add masking to their premise on the value fetched from the array, consistently with the transformation. Additionally, both step rules on array reads and writes modify their conditions for index masking, and both misspeculating rules on reads and writes tweak the side conditions under which out-of-bounds accesses are allowed.
Based on the new ideal semantics we can prove an identical set of key technical lemmas, which we elide here (see § B for details). The fact that arrays are no longer public-equivalent during misspeculation, as 5 shows, requires us to strengthen the value-based counterpart of 3 by removing that assumption, as well as an adapted version of 2: The statements of the main value SLH theorems are identical to their address SLH counterparts, substituting new translation functions for the old ones.
Theorem 6.
FvSLH enforces relative security for all IFC-well-typed programs and for all public-equivalent initial states.
Theorem 7.
SaSLH enforces SCT security for CCT programs.
VII Related Work
As already mentioned, SLH was originally proposed and implemented in LLVM [9] as a defense against Spectre v1 attacks. Patrignani and Guarnieri [20] noticed that SLH is not strong enough to achieve one of their relative security notions and proposed Strong SLH, which uses the misspeculation flag to also mask all branch conditions and all addresses of memory accesses. Zhang et al. [26] later noticed that the inputs of all variable-time instructions also have to be masked for security and implemented everything as the Ultimate SLH program transformation in the x86 backend of LLVM. They experimentally evaluated Ultimate SLH on the SPEC2017 benchmarks and reported overheads of around 150%, which is large, but still smaller than adding fences.
Taking inspiration in prior static analysis work [13, 17], both Patrignani and Guarnieri [20] and Zhang et al. [26] propose to use relative security to assess the security of their transformations, by requiring that the transformed program does not leak speculatively more than the transformed program itself leaks sequentially. While this is a sensible way to define relative security in the absence of a program transformation [13, 17, 16, 12], when applied only to the transformed program this deems secure a transformation that introduces sequential leaks, instead of removing speculative ones. Instead, we use a security definition that is more suitable for transformations and compilers, which requires that any transformed program running with speculation does not leak more than what the source program leaks sequentially. Finally, our definition only looks at 4 executions (as opposed to 8 executions [20]) and our proofs are fully mechanized in Rocq. A similar relative security definition was independently discovered by Jonathan Baumann, who verified fence-based Spectre countermeasures in the Rocq prover using hypersimulations [6].
In a separate line of research, SvSLH was proposed as an efficient way to protect cryptographic code against Spectre v1 in the Jasmin language [2]. Shivakumar et al. [22] proved in a simplified setting that an automatic SvSLH transformation (discussed in § VI) achieves speculative CT security. Yet for the best efficiency, Jasmin programmers may have to reorganize their code and manually insert a minimal number SLH protections, with a static analysis just checking that the resulting code is secure [23]. In more recent work, Arranz Olmos et al. [3] showed how speculative CT security can be preserved by compilation in a simplified model of the Jasmin compiler, which they formalized in Rocq. These preservation proofs seem easily extensible to the stronger relative security definition we use in this paper, yet providing security guarantees to non-CCT code does not seem a goal for the Jasmin language, which is specifically targeted at cryptographic implementations.
The simple and abstract speculative semantics we use in § IV is taken from the paper of Shivakumar et al. [22], who credit prior work by Cauligi et al. [11] and Barthe et al. [5] for the idea of a “forwards” semantics that takes attacker directions and does no rollbacks. This style of speculative semantics seems more suitable for higher-level languages, for which the concept of a “speculation window” that triggers rollbacks would require exposing too many low-level details about the compilation and target architecture, if at all possible to accommodate. Barthe et al. [5] also prove that in their setting any speculative CT attack against a semantics with rollbacks also exist in their “forwards” semantics with directions.
VIII Conclusion and Future Work
In this paper we introduced FSLH, a flexible variant of SLH that brings together the benefits of both Selective SLH and Ultimate SLH. We provide formal proofs in the Rocq prover that FSLH satisfies a strong relative security property, ensuring that the hardened program can only leak speculatively as much as the original source program leaks sequentially.
As a next step we want to move from the simple IFC type system of § IV-C to a flow-sensitive IFC analysis that accepts arbitrary source programs. We already implemented this in Rocq and property-based tested relative security for it using QuickChick. Formally proving this extension secure is, however, more challenging and left for future work.
On the practical side we would like to implement FSLH in LLVM and experimentally evaluate the reduction in overhead with respect to Ultimate SLH. This raises significant engineering challenges though, since it requires adding a flow-sensitive IFC analysis to LLVM, keeping track of which program inputs are secret and which ones not throughout the whole compiler chain, and making sure that the defenses we add are not removed by subsequent compiler passes. Finally, while Rocq proofs for the whole of LLVM would be way beyond what’s reasonable in terms of effort, we believe that property-based testing of relative security could be a pragmatic compromise for validating end-to-end security in practice.
Acknowledgements
We are grateful to Gilles Barthe and Santiago Arranz Olmos for the insightful discussions. This work was in part supported by the Deutsche Forschungsgemeinschaft (DFG) as part of the Excellence Strategy of the German Federal and State Governments – EXC 2092 CASA - 390781972.
Appendix A Sequential Semantics for AWhile
Figure 10 shows the standard sequential semantics for AWhile as presented in § IV-A. This figure can be derived from Figure 4 by erasing the highlighted components.
Appendix B Full Theorem Statements for FvSLH
As we point out in § VI, the statements and high-level proofs for FvSLH are very similar to their corresponding address-based variants. This starts with the relation between Flexible SLH and the other protection schemes.
Theorem 8 (Connection between FvSLH and SvSLH).
Given a set of public variables P, for any CCT program
Theorem 9 (Connection between FvSLH and USLH).
If all variables of a program (both scalars and arrays) are considered secret, i.e., , then for any program
Proof.
The assumption of public-equivalence of array states that holds for aSLH no longer applies. This change requires an adapted version of 2:
Lemma 5 (Noninterference of ).
Proof sketch.
By induction on the first evaluation judgment and inversion on the second one. ∎
This lemma is used to prove the adaptation of 3 to the vSLH setting, whose statement also removes the assumption on the public-equivalence of arrays. The same lemma is used to prove the main theorem:
Theorem 6 (FvSLH enforces relative security).
Proof.
Same structure as the proof of Theorem 3. ∎